# NbBench: Benchmarking Language Models for Comprehensive Nanobody Tasks

### v1

@@ 3 Benchmark Tasks
NbBench consists of 8 tasks across 10 datasets grouped into three categories: structure and region understanding, binding prediction, and biophysical property prediction. These tasks cover token-level and sequence-level classification, regression, and sequence generation with an overview provided in Table 1.
@@ 3.2 Binding Prediction
Antigen-Nanobody Binding Prediction (AgNbBind) predicts whether a nanobody can bind to a given antigen. This is formulated as a binary classification task. We include three separate datasets, each corresponding to a different antigen: (1) SARS-CoV-2 tsuruta2024sars and (2) human IL-6 (hIL6) tsuruta2023avida .
@@ 3.4 Data Splitting Strategy
Task #Train/Val/Test Type Granularity Metrics Source Structure and Region Understanding VRCls 13,703/3,426/2,847 Multi-class Cls Token Accuracy, Precision, Recall, F1 INDI deszynski2022indi CDRInf 13,703/3,426/2,847 Multi-class Cls Token Exact Match, BLOSUM62 Similarity INDI deszynski2022indi Binding Prediction SARS-CoV-2 27,606/6,706/38,958 Binary Cls Sequence AUROC, AUPRC, Accuracy, Precision, Recall, F1 AVIDa-SARS-CoV-2 tsuruta2024sars hIL6 13,020/11,637/449,234 Binary Cls Sequence AUROC, AUPRC, Accuracy, Precision, Recall, F1 AVIDa-hIL6 tsuruta2023avida Paratope 851/139/240 Binary Cls Token AUROC, AUPRC, Accuracy, Precision, Recall, F1 SAbDab wilton2018sdab Biophysical and Functional Property Prediction Thermo-seq 522/92/147 Regression Sequence Spearman, $R^{2}$, RMSE, MAE NbThermo/NanoMelt valdes2023nbthermo ; ramon2025prediction Thermo-tm 396/57/114 Regression Sequence Spearman, $R^{2}$, RMSE, MAE NbThermo/NanoMelt ramon2025prediction PolyRx 101,854/14,613/25,007 Binary Cls Sequence AUROC, AUPRC, Accuracy, Precision, Recall, F1 Harvey/GP-nano harvey2022silico ; zhou2024gp NbType 12,647/1,960/3,557 Multi-class Cls Sequence Accuracy, Precision, Recall, F1 PLAbDab-nano/AbNativ gordon2025plabdab ; ramon2024assessing Affinity-seq 8,888/1,302/2,547 Regression Sequence Spearman, $R^{2}$, RMSE, MAE MOQA tuvcs2024extensive Affinity-score 8,915/1,274/2,548 Regression Sequence Spearman, $R^{2}$, RMSE, MAE MOQA tuvcs2024extensive
@@ 6 Conclusions and Limitations
In this work, we introduce NbBench, the first comprehensive benchmark suite for nanobody modeling. It spans eight biologically meaningful tasks—grouped into structure understanding, antigen binding, and biophysical property prediction—and covers ten rigorously curated datasets. To ensure fair and interpretable evaluation, we systematically assess eleven pretrained language models from the protein, antibody, and nanobody domains.
@@ A.5 Full Results
Binding Prediction. We report model performance on three classification tasks: binding to SARS-CoV-2, binding to human IL-6 (hIL6), and paratope prediction. Each task is evaluated using AUROC, AUPRC, accuracy, precision, recall, and F1 score. Table 9 provides a comprehensive overview of results across these tasks. These metrics help evaluate how well models capture structural determinants of binding and functional interfaces.
@@
Table 8: Performance on SARS-CoV-2, hIL6 binding prediction, and paratope prediction tasks. Each task reports AUROC, AUPRC, Accuracy (Acc), Precision (Prec), Recall (Rec), and F1. The mean and standard deviation are reported for three different seeds.
@@
Model SARS-2 Binding hIL6 Binding Paratope Prediction AUROC AUPRC Accu Prec Rec F1 AUROC AUPRC Acc Prec Rec F1 AUROC AUPRC Acc Prec Rec F1 ProtBert 0.852 (0.004) 0.785 (0.008) 0.847 (0.005) 0.845 (0.031) 0.565 (0.020) 0.676 (0.011) 0.878 (0.011) 0.578 (0.031) 0.973 (0.002) 0.888 (0.052) 0.373 (0.076) 0.519 (0.064) 0.916 (0.004) 0.693 (0.008) 0.896 (0.000) 0.809 (0.001) 0.817 (0.001) 0.813 (0.000) ESM-2 (150M) 0.834 (0.010) 0.758 (0.012) 0.831 (0.005) 0.823 (0.023) 0.519 (0.021) 0.636 (0.015) 0.850 (0.011) 0.586 (0.026) 0.975 (0.001) 0.948 (0.001) 0.380 (0.022) 0.540 (0.023) 0.922 (0.003) 0.714 (0.002) 0.901 (0.001) 0.820 (0.012) 0.819 (0.004) 0.819 (0.003) ESM-2 (650M) 0.850 (0.002) 0.793 (0.003) 0.847 (0.003) 0.808 (0.037) 0.606 (0.033) 0.692 (0.008) 0.857 (0.005) 0.550 (0.011) 0.964 (0.002) 0.971 (0.003) 0.088 (0.053) 0.159 (0.086) 0.928 (0.000) 0.718 (0.000) 0.900 (0.000) 0.818 (0.000) 0.820 (0.000) 0.819 (0.000) AbLang-H 0.884 (0.000) 0.850 (0.000) 0.889 (0.000) 0.927 (0.000) 0.660 (0.000) 0.771 (0.000) 0.925 (0.000) 0.737 (0.000) 0.979 (0.000) 0.937 (0.000) 0.512 (0.000) 0.662 (0.000) 0.933 (0.000) 0.756 (0.000) 0.902 (0.000) 0.821 (0.000) 0.822 (0.000) 0.821 (0.000) AbLang-L 0.811 (0.018) 0.737 (0.033) 0.824 (0.013) 0.868 (0.052) 0.448 (0.019) 0.591 (0.029) 0.914 (0.011) 0.672 (0.016) 0.978 (0.002) 0.944 (0.012) 0.482 (0.003) 0.639 (0.002) 0.933 (0.000) 0.742 (0.009) 0.900 (0.000) 0.821 (0.001) 0.811 (0.018) 0.816 (0.010) AntiBERTy 0.810 (0.004) 0.719 (0.007) 0.818 (0.004) 0.852 (0.004) 0.436 (0.035) 0.576 (0.022) 0.871 (0.003) 0.565 (0.009) 0.977 (0.002) 0.932 (0.006) 0.444 (0.010) 0.602 (0.007) 0.932 (0.002) 0.739 (0.004) 0.902 (0.001) 0.821 (0.003) 0.821 (0.003) 0.821 (0.002) AntiBERTa2 0.824 (0.001) 0.760 (0.001) 0.834 (0.000) 0.769 (0.001) 0.596 (0.000) 0.672 (0.000) 0.887 (0.006) 0.615 (0.020) 0.976 (0.001) 0.926 (0.013) 0.445 (0.032) 0.597 (0.025) 0.930 (0.001) 0.746 (0.001) 0.903 (0.002) 0.825 (0.001) 0.816 (0.020) 0.821 (0.010) AntiBERTa2-CSSP 0.845 (0.011) 0.793 (0.013) 0.850 (0.011) 0.831 (0.021) 0.591 (0.043) 0.690 (0.028) 0.916 (0.011) 0.725 (0.015) 0.981 (0.001) 0.939 (0.000) 0.548 (0.020) 0.696 (0.018) 0.939 (0.013) 0.765 (0.004) 0.906 (0.002) 0.831 (0.001) 0.822 (0.001) 0.826 (0.001) IgBert 0.817 (0.003) 0.743 (0.008) 0.817 (0.014) 0.909 (0.045) 0.399 (0.082) 0.549 (0.068) 0.821 (0.003) 0.430 (0.020) 0.969 (0.000) 0.929 (0.022) 0.247 (0.005) 0.389 (0.006) 0.921 (0.001) 0.687 (0.002) 0.896 (0.001) 0.808 (0.002) 0.820 (0.006) 0.814 (0.002) NanoBERT 0.876 (0.017) 0.829 (0.029) 0.874 (0.019) 0.873 (0.039) 0.650 (0.039) 0.745 (0.040) 0.904 (0.015) 0.687 (0.021) 0.978 (0.001) 0.948 (0.009) 0.481 (0.007) 0.638 (0.005) 0.928 (0.001) 0.719 (0.002) 0.897 (0.001) 0.812 (0.004) 0.810 (0.004) 0.811 (0.001) VHHBERT 0.800 (0.010) 0.712 (0.007) 0.811 (0.007) 0.811 (0.082) 0.454 (0.101) 0.573 (0.061) 0.873 (0.000) 0.593 (0.000) 0.976 (0.000) 0.955 (0.000) 0.426 (0.000) 0.589 (0.000) 0.921 (0.002) 0.695 (0.007) 0.890 (0.001) 0.802 (0.002) 0.794 (0.004) 0.798 (0.002)
@@
Table 9: Performance on Thermo-seq, Thermo-tm, Affinity-seq and Affinity-score tasks. Each task reports Spearman coefficient, $R^{2}$, RMSE and MAE.
@@
Model Thermo-seq Thermo-tm Affinity-seq Affinity-score Spearman $R^{2}$ RMSE MAE Spearman $R^{2}$ RMSE MAE Spearman $R^{2}$ RMSE MAE Spearman $R^{2}$ RMSE MAE ProtBert 0.040 (0.001) 0.003 (0.000) 20.689 (0.048) 18.767 (0.052) 0.128 (0.002) 0.013 (0.000) 24.797 (0.054) 22.886 (0.107) 0.163 (0.008) 0.044 (0.006) 4.003 (0.025) 3.358 (0.029) 0.084 (0.015) 0.030 (0.008) 1.787 (0.006) 1.296 (0.119) ESM-2 (150M) 0.389 (0.000) 0.273 (0.000) 9.946 (0.000) 7.771 (0.000) 0.301 (0.003) 0.113 (0.001) 9.255 (0.025) 7.810 (0.025) 0.170 (0.000) 0.073 (0.000) 3.935 (0.000) 3.319 (0.000) 0.063 (0.000) 0.037 (0.000) 1.842 (0.000) 1.066 (0.000) ESM-2 (650M) 0.375 (0.010) 0.154 (0.005) 10.099 (0.050) 7.683 (0.026) 0.315 (0.000) 0.096 (0.000) 9.178 (0.000) 7.842 (0.000) 0.183 (0.003) 0.082 (0.001) 3.464 (0.006) 2.779 (0.003) 0.089 (0.001) 0.052 (0.006) 1.855 (0.036) 1.267 (0.236) AbLang-H 0.566 (0.005) 0.413 (0.012) 10.482 (0.227) 8.239 (0.146) 0.533 (0.004) 0.272 (0.006) 9.916 (0.110) 7.936 (0.107) 0.151 (0.000) 0.025 (0.000) 5.031 (0.000) 4.537 (0.000) 0.106 (0.000) 0.063 (0.000) 1.898 (0.000) 1.608 (0.000) AbLang-L 0.585 (0.001) 0.412 (0.002) 9.161 (0.171) 7.173 (0.118) 0.502 (0.002) 0.264 (0.003) 8.855 (0.105) 7.143 (0.054) 0.173 (0.010) 0.053 (0.008) 4.073 (0.722) 3.316 (0.809) 0.128 (0.004) 0.058 (0.007) 2.082 (0.125) 1.821 (0.153) AntiBERTy 0.474 (0.006) 0.345 (0.009) 9.187 (0.204) 7.050 (0.119) 0.385 (0.007) 0.140 (0.003) 9.640 (0.069) 7.888 (0.063) 0.167 (0.017) 0.074 (0.017) 3.532 (0.456) 3.053 (0.544) 0.119 (0.013) 0.072 (0.003) 1.915 (0.106) 1.408 (0.194) AntiBERTa2 0.587 (0.002) 0.451 (0.000) 8.493 (0.001) 6.533 (0.000) 0.585 (0.003) 0.343 (0.002) 8.615 (0.142) 6.723 (0.145) 0.162 (0.025) 0.041 (0.007) 3.534 (0.731) 2.840 (0.844) 0.088 (0.001) 0.053 (0.009) 2.150 (0.092) 1.630 (0.409) AntiBERTa2-CSSP 0.575 (0.010) 0.459 (0.010) 8.808 (0.151) 6.708 (0.097) 0.593 (0.004) 0.361 (0.004) 8.469 (0.109) 6.839 (0.083) 0.149 (0.003) 0.067 (0.006) 3.921 (0.940) 3.307 (1.052) 0.085 (0.011) 0.066 (0.014) 2.057 (0.113) 1.418 (0.450) IgBert 0.351 (0.178) 0.156 (0.098) 10.028 (0.344) 7.873 (0.168) 0.344 (0.176) 0.144 (0.111) 9.109 (0.358) 7.544 (0.337) 0.184 (0.006) 0.082 (0.013) 3.369 (0.582) 2.670 (0.666) 0.090 (0.037) 0.065 (0.001) 1.908 (0.236) 1.285 (0.185) NanoBERT 0.440 (0.006) 0.241 (0.009) 9.590 (0.073) 7.380 (0.057) 0.298 (0.003) 0.053 (0.001) 10.011 (0.037) 7.949 (0.013) 0.061 (0.000) 0.006 (0.000) 3.518 (0.000) 2.811 (0.000) 0.118 (0.000) 0.022 (0.000) 1.895 (0.000) 1.575 (0.000) VHHBERT 0.484 (0.007) 0.390 (0.004) 9.318 (0.053) 7.540 (0.021) 0.532 (0.000) 0.310 (0.000) 8.701 (0.000) 6.883 (0.000) 0.145 (0.015) 0.044 (0.019) 3.110 (0.152) 2.248 (0.178) 0.115 (0.006) 0.070 (0.003) 2.092 (0.130) 1.667 (0.402)

### v2

@@ 3 Benchmark Tasks
NbBench consists of 8 tasks across 9 datasets grouped into three categories: structure and region understanding, binding prediction, and biophysical property prediction. These tasks cover token-level and sequence-level classification, regression, and sequence generation with an overview provided in Table 1.
@@ 3.2 Binding Prediction
Antigen-Nanobody Binding Prediction (AgNbBind) predicts whether a nanobody can bind to a given antigen. This is formulated as a binary classification task. We include two separate datasets, each corresponding to a different antigen: (1) SARS-CoV-2 tsuruta2024sars and (2) human IL-6 (hIL6) tsuruta2023avida .
@@ 3.4 Data Splitting Strategy
Task #Train/Val/Test Type Granularity Metrics Source Structure and Region Understanding VRCls 13,703/3,426/2,847 Multi-class Cls Token Accuracy, Precision, Recall, F1 INDI deszynski2022indi CDRInf 13,703/3,426/2,847 Multi-class Cls Token Exact Match, BLOSUM62 Similarity INDI deszynski2022indi Binding Prediction SARS-CoV-2 27,606/6,706/38,958 Binary Cls Sequence AUROC, AUPRC, Accuracy, Precision, Recall, F1 AVIDa-SARS-CoV-2 tsuruta2024sars hIL6 13,020/11,637/449,234 Binary Cls Sequence AUROC, AUPRC, Accuracy, Precision, Recall, F1 AVIDa-hIL6 tsuruta2023avida Paratope 851/139/240 Binary Cls Token AUROC, AUPRC, Accuracy, Precision, Recall, F1 SAbDab wilton2018sdab Biophysical and Functional Property Prediction Thermo-seq 522/92/147 Regression Sequence Spearman, $R^{2}$, RMSE, MAE NbThermo/NanoMelt/TEMPRO valdes2023nbthermo ; ramon2025prediction ; alvarez2024tempro Thermo-tm 396/57/114 Regression Sequence Spearman, $R^{2}$, RMSE, MAE NbThermo/NanoMelt/TEMPRO valdes2023nbthermo ; ramon2025prediction ; alvarez2024tempro PolyRx 101,854/14,613/25,007 Binary Cls Sequence AUROC, AUPRC, Accuracy, Precision, Recall, F1 Harvey/GP-nano harvey2022silico ; zhou2024gp NbType 12,647/1,960/3,557 Multi-class Cls Sequence Accuracy, Precision, Recall, F1 PLAbDab-nano/AbNativ gordon2025plabdab ; ramon2024assessing Affinity-seq 8,888/1,302/2,547 Regression Sequence Spearman, $R^{2}$, RMSE, MAE MOQA tuvcs2024extensive Affinity-score 8,915/1,274/2,548 Regression Sequence Spearman, $R^{2}$, RMSE, MAE MOQA tuvcs2024extensive
@@ 5.2 Benchmark Results
Model VRCls CDRInf SARS-2 hIL6 Paratope Thermo-seq Thermo-tm PolyRx NbType Affinity-seq Affinity-score
Metric Accuracy BR AUROC AUROC AUROC Spearman Spearman AUROC Accuracy Spearman Spearman
Protein Language Models
[…]
Antibody Language Models
[…]
Nanobody Language Models
[…]
@@ 6 Conclusions and Limitations
In this work, we introduce NbBench, the first comprehensive benchmark suite for nanobody modeling. It spans eight biologically meaningful tasks—grouped into structure understanding, antigen binding, and biophysical property prediction—and covers nine rigorously curated datasets. To ensure fair and interpretable evaluation, we systematically assess eleven pretrained language models from the protein, antibody, and nanobody domains.
@@ A.5 Full Results
Binding Prediction. We report model performance on three classification tasks: binding to SARS-CoV-2, binding to human IL-6 (hIL6), and paratope prediction. Each task is evaluated using AUROC, AUPRC, accuracy, precision, recall, and F1 score. Table LABEL:tab:binding-metrics provides a comprehensive overview of results across these tasks. These metrics help evaluate how well models capture structural determinants of binding and functional interfaces.
@@
Table 8: Performance of various models across three tasks: SARS-CoV-2 Binding Prediction, hIL6 Binding Prediction, and Paratope Prediction. Each metric reports the mean and standard deviation over three seeds.
@@
Task Model AUROC AUPRC Acc Prec Rec F1 SARS-CoV-2 ProtBert 0.852 (0.004) 0.785 (0.008) 0.847 (0.005) 0.845 (0.031) 0.565 (0.020) 0.676 (0.011) ESM-2 (150M) 0.834 (0.010) 0.758 (0.012) 0.831 (0.005) 0.823 (0.023) 0.519 (0.021) 0.636 (0.015) ESM-2 (650M) 0.850 (0.002) 0.793 (0.003) 0.847 (0.003) 0.808 (0.037) 0.606 (0.033) 0.692 (0.008) AbLang-H 0.884 (0.000) 0.850 (0.000) 0.889 (0.000) 0.927 (0.000) 0.660 (0.000) 0.771 (0.000) AbLang-L 0.811 (0.018) 0.737 (0.033) 0.824 (0.013) 0.868 (0.052) 0.448 (0.019) 0.591 (0.029) AntiBERTy 0.810 (0.004) 0.719 (0.007) 0.818 (0.004) 0.852 (0.004) 0.436 (0.035) 0.576 (0.022) AntiBERTa2 0.824 (0.001) 0.760 (0.001) 0.834 (0.000) 0.769 (0.001) 0.596 (0.000) 0.672 (0.000) AntiBERTa2-CSSP 0.845 (0.011) 0.793 (0.013) 0.850 (0.011) 0.831 (0.021) 0.591 (0.043) 0.690 (0.028) IgBert 0.817 (0.003) 0.743 (0.008) 0.817 (0.014) 0.909 (0.045) 0.399 (0.082) 0.549 (0.068) NanoBERT 0.876 (0.017) 0.829 (0.029) 0.874 (0.019) 0.873 (0.039) 0.650 (0.039) 0.745 (0.040) VHHBERT 0.800 (0.010) 0.712 (0.007) 0.811 (0.007) 0.811 (0.082) 0.454 (0.101) 0.573 (0.061) hIL6 ProtBert 0.878 (0.011) 0.578 (0.031) 0.973 (0.002) 0.888 (0.052) 0.373 (0.076) 0.519 (0.064) ESM-2 (150M) 0.850 (0.011) 0.586 (0.026) 0.975 (0.001) 0.948 (0.001) 0.380 (0.022) 0.540 (0.023) ESM-2 (650M) 0.857 (0.005) 0.550 (0.011) 0.964 (0.002) 0.971 (0.003) 0.088 (0.053) 0.159 (0.086) AbLang-H 0.925 (0.000) 0.737 (0.000) 0.979 (0.000) 0.937 (0.000) 0.512 (0.000) 0.662 (0.000) AbLang-L 0.914 (0.011) 0.672 (0.016) 0.978 (0.002) 0.944 (0.012) 0.482 (0.003) 0.639 (0.002) AntiBERTy 0.871 (0.003) 0.565 (0.009) 0.977 (0.002) 0.932 (0.006) 0.444 (0.010) 0.602 (0.007) AntiBERTa2 0.887 (0.006) 0.615 (0.020) 0.976 (0.001) 0.926 (0.013) 0.445 (0.032) 0.597 (0.025) AntiBERTa2-CSSP 0.916 (0.011) 0.725 (0.015) 0.981 (0.001) 0.939 (0.000) 0.548 (0.020) 0.696 (0.018) IgBert 0.821 (0.003) 0.430 (0.020) 0.969 (0.000) 0.929 (0.022) 0.247 (0.005) 0.389 (0.006) NanoBERT 0.904 (0.015) 0.687 (0.021) 0.978 (0.001) 0.948 (0.009) 0.481 (0.007) 0.638 (0.005) VHHBERT 0.873 (0.000) 0.593 (0.000) 0.976 (0.000) 0.955 (0.000) 0.426 (0.000) 0.589 (0.000) Paratope ProtBert 0.916 (0.004) 0.693 (0.008) 0.896 (0.000) 0.809 (0.001) 0.817 (0.001) 0.813 (0.000) ESM-2 (150M) 0.922 (0.003) 0.714 (0.002) 0.901 (0.001) 0.820 (0.012) 0.819 (0.004) 0.819 (0.003) ESM-2 (650M) 0.928 (0.000) 0.718 (0.000) 0.900 (0.000) 0.818 (0.000) 0.820 (0.000) 0.819 (0.000) AbLang-H 0.933 (0.000) 0.756 (0.000) 0.902 (0.000) 0.821 (0.000) 0.822 (0.000) 0.821 (0.000) AbLang-L 0.933 (0.000) 0.742 (0.009) 0.900 (0.000) 0.821 (0.001) 0.811 (0.018) 0.816 (0.010) AntiBERTy 0.932 (0.002) 0.739 (0.004) 0.902 (0.001) 0.821 (0.003) 0.821 (0.003) 0.821 (0.002) AntiBERTa2 0.930 (0.001) 0.746 (0.001) 0.903 (0.002) 0.825 (0.001) 0.816 (0.020) 0.821 (0.010) AntiBERTa2-CSSP 0.939 (0.013) 0.765 (0.004) 0.906 (0.002) 0.831 (0.001) 0.822 (0.001) 0.826 (0.001) IgBert 0.921 (0.001) 0.687 (0.002) 0.896 (0.001) 0.808 (0.002) 0.820 (0.006) 0.814 (0.002) NanoBERT 0.928 (0.001) 0.719 (0.002) 0.897 (0.001) 0.812 (0.004) 0.810 (0.004) 0.811 (0.001) VHHBERT 0.921 (0.002) 0.695 (0.007) 0.890 (0.001) 0.802 (0.002) 0.794 (0.004) 0.798 (0.002)
@@
Table 9: Performance on regression tasks (Thermo-seq, Thermo-tm, Affinity-seq and Affinity-score). Each task reports Spearman coefficient, $R^{2}$, RMSE and MAE (mean $\pm$ std).
@@
Model Thermo-seq Thermo-tm Spearman $R^{2}$ RMSE MAE Spearman $R^{2}$ RMSE MAE ProtBert 0.040 (0.001) 0.003 (0.000) 20.689 (0.048) 18.767 (0.052) 0.128 (0.002) 0.013 (0.000) 24.797 (0.054) 22.886 (0.107) ESM-2 (150M) 0.389 (0.000) 0.273 (0.000) 9.946 (0.000) 7.771 (0.000) 0.301 (0.003) 0.113 (0.001) 9.255 (0.025) 7.810 (0.025) ESM-2 (650M) 0.375 (0.010) 0.154 (0.005) 10.099 (0.050) 7.683 (0.026) 0.315 (0.000) 0.096 (0.000) 9.178 (0.000) 7.842 (0.000) AbLang-H 0.566 (0.005) 0.413 (0.012) 10.482 (0.227) 8.239 (0.146) 0.533 (0.004) 0.272 (0.006) 9.916 (0.110) 7.936 (0.107) AbLang-L 0.585 (0.001) 0.412 (0.002) 9.161 (0.171) 7.173 (0.118) 0.502 (0.002) 0.264 (0.003) 8.855 (0.105) 7.143 (0.054) AntiBERTy 0.474 (0.006) 0.345 (0.009) 9.187 (0.204) 7.050 (0.119) 0.385 (0.007) 0.140 (0.003) 9.640 (0.069) 7.888 (0.063) AntiBERTa2 0.587 (0.002) 0.451 (0.000) 8.493 (0.001) 6.533 (0.000) 0.585 (0.003) 0.343 (0.002) 8.615 (0.142) 6.723 (0.145) AntiBERTa2-CSSP 0.575 (0.010) 0.459 (0.010) 8.808 (0.151) 6.708 (0.097) 0.593 (0.004) 0.361 (0.004) 8.469 (0.109) 6.839 (0.083) IgBert 0.351 (0.178) 0.156 (0.098) 10.028 (0.344) 7.873 (0.168) 0.344 (0.176) 0.144 (0.111) 9.109 (0.358) 7.544 (0.337) NanoBERT 0.440 (0.006) 0.241 (0.009) 9.590 (0.073) 7.380 (0.057) 0.298 (0.003) 0.053 (0.001) 10.011 (0.037) 7.949 (0.013) VHHBERT 0.484 (0.007) 0.390 (0.004) 9.318 (0.053) 7.540 (0.021) 0.532 (0.000) 0.310 (0.000) 8.701 (0.000) 6.883 (0.000)
Model Affinity-seq Affinity-score Spearman $R^{2}$ RMSE MAE Spearman $R^{2}$ RMSE MAE ProtBert 0.163 (0.008) 0.044 (0.006) 4.003 (0.025) 3.358 (0.029) 0.084 (0.015) 0.030 (0.008) 1.787 (0.006) 1.296 (0.119) ESM-2 (150M) 0.170 (0.000) 0.073 (0.000) 3.935 (0.000) 3.319 (0.000) 0.063 (0.000) 0.037 (0.000) 1.842 (0.000) 1.066 (0.000) ESM-2 (650M) 0.183 (0.003) 0.082 (0.001) 3.464 (0.006) 2.779 (0.003) 0.089 (0.001) 0.052 (0.006) 1.855 (0.036) 1.267 (0.236) AbLang-H 0.151 (0.000) 0.025 (0.000) 5.031 (0.000) 4.537 (0.000) 0.106 (0.000) 0.063 (0.000) 1.898 (0.000) 1.608 (0.000) AbLang-L 0.173 (0.010) 0.053 (0.008) 4.073 (0.722) 3.316 (0.809) 0.128 (0.004) 0.058 (0.007) 2.082 (0.125) 1.821 (0.153) AntiBERTy 0.167 (0.017) 0.074 (0.017) 3.532 (0.456) 3.053 (0.544) 0.119 (0.013) 0.072 (0.003) 1.915 (0.106) 1.408 (0.194) AntiBERTa2 0.162 (0.025) 0.041 (0.007) 3.534 (0.731) 2.840 (0.844) 0.088 (0.001) 0.053 (0.009) 2.150 (0.092) 1.630 (0.409) AntiBERTa2-CSSP 0.149 (0.003) 0.067 (0.006) 3.921 (0.940) 3.307 (1.052) 0.085 (0.011) 0.066 (0.014) 2.057 (0.113) 1.418 (0.450) IgBert 0.184 (0.006) 0.082 (0.013) 3.369 (0.582) 2.670 (0.666) 0.090 (0.037) 0.065 (0.001) 1.908 (0.236) 1.285 (0.185) NanoBERT 0.061 (0.000) 0.006 (0.000) 3.518 (0.000) 2.811 (0.000) 0.118 (0.000) 0.022 (0.000) 1.895 (0.000) 1.575 (0.000) VHHBERT 0.145 (0.015) 0.044 (0.019) 3.110 (0.152) 2.248 (0.178) 0.115 (0.006) 0.070 (0.003) 2.092 (0.130) 1.667 (0.402)